Problems:
The usual problems tracking/guiding in the north., but also around Dec= -26 .Again had some success far north at 02h +15d for a while.
No other telescope problems.
Very few SISPI problems tonight. No Interlock errors. One instance of exposure froze before readout; had to hit "Stop" (154661). Image written to disk.